# On the maximal correlation of some stochastic processes

### v2

@@ 4 Lévy processes and stable distributions
In this subsection, we prove Theorem 1.2. The proof crucially relies on the Lévy-Itô decomposition and the martingale representation theorem for Lévy processes. We will briefly present these classical results. We focus on two-dimensional Lévy processes. Without loss of generality, we may assume that the Lévy process is canonical. Take the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ as the sample space $\Omega$, where $D_{\mathbb{R}^{2}}[0,\infty)$ is the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology and the corresponding Borel $\sigma$-field. For $\omega=(\omega_{t})_{t\geq 0}$, define $(X_{t},Y_{t})=\omega_{t}$ for $t\geq 0$. Let $P$ be a probability measure on $\Omega$ such that $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process. Let
@@
be the jump of $(X_{t})_{t\geq 0}$ at time $t$. Similarly, we define $\Delta Y_{t}$. Then, the set $\{t\geq 0:\Delta X_{t}\neq 0\text{ or }\Delta Y_{t}\neq 0\}$ of jumping times is countable.
@@
If $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process, then there exists $b=(b^{X},b^{Y})\in\mathbb{R}^{2}$, a two-dimensional Brownian motion $(B^{X}_{t},B^{Y}_{t})_{t\geq 0}$ with the covariance $\Sigma$ and a Poisson random measure $N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ with the intensity measure $\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y)$ such that for $t\geq 0$,
@@
where $\delta_{z_{0}}(\mathrm{d}z)$ is the Dirac measure at $z_{0}\in[0,\infty)\times(\mathbb{R}^{2}\setminus\{0\})$.
We refer to [App09, Theorem 2.4.16] for a proof. Note that the random measure $N^{X}(\mathrm{d}t,\mathrm{d}x)$ is a measurable function of $(X_{t})_{t\geq 0}$. Here, the space of $\sigma$-finite measures is endowed with the weak topology. Moreover, almost surely, $t\mapsto\int_{0}^{t}\int_{|x|\geq 1}xN^{X}(\mathrm{d}s,\mathrm{d}x)$ is a random element in $D[0,\infty)$. Hence, the process $t\mapsto\int_{0}^{t}\int_{|x|\geq 1}xN^{X}(\mathrm{d}s,\mathrm{d}x)$ in $D[0,\infty)$ can be realized as a measurable function of $(X_{t})_{t\geq 0}$ after a modification up to a $P$-null set. For $n\geq 1$, define
$Z^{(n)}_{t}=\int_{0}^{t}\int_{1/n<|x|<1}x\widetilde{N}^{X}(\mathrm{d}s,\mathrm{d}x).$
Then, each $(Z^{(n)}_{t})_{t\geq 0}$ is a càdlàg martingale, and can be viewed as measurable function of $(X_{t})_{t\geq 0}$ with values in $D[0,\infty)$. Let
$Z_{t}=\int_{0}^{t}\int_{|x|<1}x\widetilde{N}^{X}(\mathrm{d}s,\mathrm{d}x).$
The process $(Z_{t})_{t\geq 0}$ is viewed as a càdlàg martingale, which is the limit of $(Z^{(n)}_{t})_{t\geq 0}$. To be more precise, it is proved that for fixed $t\geq 0$, $Z^{(n)}_{t}$ converges to $Z_{t}$ in $L^{2}$. Hence, by using Doob’s maximal inequality and taking subsequences if necessary, there exists a subsequence $Z^{(n_{k})}$ such that $Z^{(n_{k})}$ converges almost surely to $Z$ in the Skorokhod space $D[0,\infty)$. Hence, after a modification up to a $P$-null set, $(Z_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$. Then, by (14), the Brownian motion $(B^{X}_{t})_{t\geq 0}$ is also a measurable function of $(X_{t})_{t\geq 0}$. Similarly, $(B^{Y}_{t})_{t\geq 0}$ and $N^{Y}(\mathrm{d}t,\mathrm{d}y)$ are measurable functions of $(Y_{t})_{t\geq 0}$.
@@
Let $F$ be a square integrable $\mathcal{F}^{X}_{T}$-measurable random variable. Then, there exist square-integrable predictable processes $\varphi^{X}$ and $\psi^{X}$ such that almost surely,
@@
Now, let’s start to prove Theorem 1.2. Since $(B^{X}_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$, $B^{X}_{1}$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$. Similarly, $B^{Y}_{1}$ is a square-integrable measurable function of $(Y_{t})_{t\geq 0}$. Hence, we have that
@@
Next, by definition of $\mathrm{Op}(\nu)$, for each small $\varepsilon>0$, there exists measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$, $\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)>0$ and that
@@
Then, $F$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$ and $G$ is a square integrable measurable function of $(Y_{t})_{t\geq 0}$. Moreover, we have that $E(F)=E(G)=0$ and that
@@
To prove the reverse inequality $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\max(|\rho|,\mathrm{Op}(\nu))$, we fix $T>0$ and non-degenerate square-integrable $F=F((X_{t})_{0\leq t\leq T})$ and $G=G((Y_{t})_{0\leq t\leq T})$, where $F(\cdot)$ and $G(\cdot)$ are measurable functions. By [Bil99, Theorem 12.5], the Borel-$\sigma$-field $\mathcal{D}$ of the Skorokhod space $D[0,T]$ is equal to the sigma field generated by finite-dimensional cylinders. Hence, we have that $F\in\sigma(X_{t},0\leq t\leq T)\subset\mathcal{F}^{X}_{T}$ and $G\in\sigma(Y_{t},0\leq t\leq T)\subset\mathcal{F}^{Y}_{T}$. By Theorem 4.2, we find square-integrable predictable processes $\varphi^{X}$, $\psi^{X}$, $\varphi^{Y}$ and $\psi^{Y}$ such that $F=E(F)+F_{B}+F_{N}$ and $G=E(G)+G_{B}+G_{N}$ with probability $1$, where
@@
We only provide the proof for $E(F_{B}G_{N})=0$ as the other equations can be proved in a similar way. We define a martingale-valued measure $M$ as follows: for $t\geq 0$ and $A\subset\mathbb{R}^{2}$, we define
@@
Let $\mu(t,A)=E(M(t,A)^{2})$. Then, by independence between $\widetilde{N}$ and $B^{X}$, we get that
@@
Let $f(t,x,y)=1_{x=y=0}\varphi^{X}(t)$ and $g(t,x,y)=1_{y\neq 0}\psi^{Y}(t,y)$. Then, $F_{B}$ equals to the stochastic integration $\int_{0}^{T}f(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ and $G_{N}$ equals to the stochastic integration $\int_{0}^{T}g(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$. For the general theory of stochastic integration against certain type of martingale-valued measure, we refer to [App09, Chapter 4]. By Itô’s isometry for stochastic integrals (see e.g. [App09, Theorem 4.2.3 and Exercise 4.2.4]), we have that
@@
[…]
@@
By Cauchy-Schwarz inequality, we obtain that
@@
Integrating both sides of (21) against $\mathrm{d}t$ and using Cauchy-Schwarz inequality, we prove (20). Finally, by (18), (19), (20) and Cauchy-Schwarz inequality, we have that
@@
As $T\to\infty$, $((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})$ converges to $((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$, by Lemma 2.3 (lower-semi continuity), we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\liminf_{T\to\infty}R((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})\leq\max(|\rho|,\mathrm{Op}(\nu))$
@@
We will prove Theorem 1.3 in this subsection. Firstly, we briefly explain the ideas in the following: We have found the maximal correlation coefficients for two dimensional Lévy processes in Theorem 1.2. Then, we apply Theorem 1.2 to two dimensional $\alpha$-stable processes $(X_{t},Y_{t})_{t\geq 0}$ with $20<\alpha<$ 2. The Brownian part vanishes. Hence, it suffices to calculate $\mathrm{Op}(\nu)$. For $\alpha$-stable processes, $\nu$ takes the special form (9). When $\tau(\mathrm{d}\theta)$ has a density $\tau(\theta)$ with respect to the Lebesgue measure $\mathrm{d}\theta$, the Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$ also has a density $\nu(x,y)$. Moreover, if $\tau(-\theta)=\tau(\theta)$, we have $\nu(-x,-y)=\nu(x,y)$. In this case, we can define a homogeneous kernel $K(x,y)=\frac{\nu(x,y)}{\sqrt{\nu_{X}(x)\nu_{Y}(y)}}$ of degree $-1$, where $\nu_{X}(x)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}y$, $\nu_{Y}(y)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}x$ and $K(\lambda x,\lambda y)=|\lambda|^{-1}K(x,y)$ for all $x,y\in\mathbb{R}$ and $\lambda\neq 0$. The kernel $K(x,y)$ induces a linear operator $K$ on $L^{2}$ by $K\psi(x):=\int_{\mathbb{R}}K(x,y)\psi(y)\,\mathrm{d}y$. Then, the constant $\mathrm{Op}(\nu)$ is just the operator norm $||K||=\sup_{||\psi||_{2}>0}\frac{||K\psi||_{2}}{||\psi||_{2}}$. Here, “$\mathrm{Op}$” is short for the word “operator”. The determination of norm $||K||$ of homogenous kernel of degree $-1$ is a classical problem under the name Hilbert-Hardy inequality, see [HLP88, Theorem 319] and [Yan12, Theorem 42.8]. However, in our case, $\tau(\mathrm{d}\theta)$ is not necessarily absolutely continuous with respect to the Lebesgue measure. Moreover, even it is possible to define the kernel $K(x,y)$ in certain cases, due to the absence of the symmetry $\tau(-\theta)=\tau(\theta)$, $K(x,y)$ is not homogeneous. Instead, $K(x,y)$ is only positively homogeneous, i.e. $K(\lambda x,\lambda y)=\lambda^{-1}K(x,y)$ for $x,y\in\mathbb{R}$ and $\lambda>0$. The difference between homogeneity and positive homogeneity results in significant different expressions of $||K||$. We are not aware of existing results for $\mathrm{Op}(\nu)$. Hence, we calculate $\mathrm{Op}(\nu)$ in Lemma 4.4 by an adaptation of the argument leading to the Hilbert-Hardy inequalities. In this way, we successively find the maximal correlation coefficients $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for $\alpha$-stable processes. As $\alpha$-stable distributions are marginal distributions of $\alpha$-stable processes, we have the upper bound $R(X_{1},Y_{1})\leq R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. The reverse inequality is a consequence of the convergence of stable random walks towards stable processes, Theorem 1.1 and Lemma 2.3. By the way, we could not generalize Theorem 1.3 to general infinitely divisible distributions. Indeed, it is possible to have $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for general two dimensional Lévy processes.
@@
Then, we have that
$\mathrm{Op}(\nu)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $/0=0$ 0, where $\|\cdot\|_{2}$ means the spectral norm.
@@
using Cauchy-Schwarz inequality, we obtain that
@@
$\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{x}_{+}F_{+}+D^{x}_{-}F_{-},$ (24)
where
@@
$\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{y}_{+}G_{+}+D^{y}_{-}G_{-},$ (25)
where
@@
$A=\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}$
with the convention that $/0=0$ 0. Let $||A||_{2}$ be the spectral norm of $A$. Then by (23), (24) and (25), we have that
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\leq||A||_{2}\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}.$
Hence, $\mathrm{Op}(\nu)\leq||A||_{2}$. It remains to prove the converse inequality $\mathrm{Op}(\nu)\geq||A||_{2}$. Without loss of generality, we may assume that $||A||_{2}>0$. Equivalently, we assume that $D^{x}_{+}+D^{x}_{-}>0$ and $D^{y}_{+}+D^{y}_{-}>0$. For this purpose, for $\varepsilon>0$, we set
@@
where the positive constants $b^{\varphi}_{+}$, $b^{\varphi}_{-}$, $b^{\psi}_{+}$ and $b^{\psi}_{-}$ will be chosen later. Then, we have that
@@
$(\sqrt{D^{x}_{+}}b^{\varphi}_{+},\sqrt{D^{x}_{-}}b^{\varphi}_{-})A\begin{pmatrix}\sqrt{D^{y}_{+}}b^{\psi}_{+}\\ \sqrt{D^{y}_{-}}b^{\psi}_{-}\end{pmatrix}\\ =||A||_{2}\sqrt{(D^{x}_{+}(b^{\varphi}_{+})^{2}+D^{x}_{-}(b^{\varphi}_{-})^{2})(D^{y}_{+}(b^{\psi}_{+})^{2}+D^{y}_{-}(b^{\psi}_{-})^{2})}.$
Then, we get that
$\mathrm{Op}(\nu)\geq\lim_{\varepsilon\to 0}\frac{\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)}{\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}}=||A||_{2}.$
At present, we have found the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. Clearly, we have that
@@
On the other hand, consider the random walk $(S_{n},T_{n})_{n\geq 0}$ such that its increment $(S_{n+1}-S_{n},T_{n+1}-T_{n})$ has the same distribution as $(X_{1},Y_{1})$. Then, by Theorem 1.1, we have that

### v3

@@ 4 Two-dimensional Lévy processes and bivariate stable distributions
In this subsection, we prove Theorem 1.2. The proof relies crucially on the Lévy-Itô decomposition and the martingale representation theorem for Lévy processes. We will briefly present these classical results. We focus on two-dimensional Lévy processes. Without loss of generality, we may assume that the Lévy process is canonical. Take the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ as the sample space $\Omega$, where $D_{\mathbb{R}^{2}}[0,\infty)$ is the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology and the corresponding Borel $\sigma$-field. For $\omega=(\omega_{t})_{t\geq 0}$, define $(X_{t},Y_{t})=\omega_{t}$ for $t\geq 0$. Let $P$ be a probability measure on $\Omega$ such that $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process. Let
@@
be the jump of $(X_{t})_{t\geq 0}$ at time $t$. Similarly, we define $\Delta Y_{t}$. Then the set $\{t\geq 0:\Delta X_{t}\neq 0\text{ or }\Delta Y_{t}\neq 0\}$ of jumping times is countable.
@@
If $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process, then there exists $b=(b^{X},b^{Y})\in\mathbb{R}^{2}$, a two-dimensional Brownian motion $(B^{X}_{t},B^{Y}_{t})_{t\geq 0}$ with the covariance matrix $\Sigma$ and a Poisson random measure $N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ with the intensity measure $\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y)$ such that for $t\geq 0$,
@@
where $\delta_{z_{0}}(\mathrm{d}z)$ is the Dirac measure at $z_{0}\in[0,\infty)\times(\mathbb{R}^{2}\setminus\{(0,0)\})$. In addition, $(B^{X}_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$, and $(B^{Y}_{t})_{t\geq 0}$ is a measurable function of $(Y_{t})_{t\geq 0}$.
We refer to [App09, Theorem 2.4.16] for a proof.
@@
Let $F$ be a square-integrable $\mathcal{F}^{X}_{T}$-measurable random variable. Then there exist square-integrable predictable processes $\varphi^{X}$ and $\psi^{X}$ such that almost surely,
@@
Now, let us start to prove Theorem 1.2. Since $(B^{X}_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$, $B^{X}_{1}$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$. Similarly, $B^{Y}_{1}$ is a square-integrable measurable function of $(Y_{t})_{t\geq 0}$. Hence, we have that
@@
Next, by the definition of $\mathrm{Op}(\nu)$, for each small $\varepsilon>0$, there exist measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$, $0<\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)<\infty$ and that
@@
Then $F$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$ and $G$ is a square-integrable measurable function of $(Y_{t})_{t\geq 0}$. Moreover, we have that $E(F)=E(G)=0$ and that
@@
To prove the reverse inequality $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\max(|\rho|,\mathrm{Op}(\nu))$, we fix $T>0$ and square-integrable non-degenerate $F=F((X_{t})_{0\leq t\leq T})$ and $G=G((Y_{t})_{0\leq t\leq T})$, where $F(\cdot)$ and $G(\cdot)$ are measurable functions. By [Bil99, Theorem 12.5], the Borel $\sigma$-field $\mathcal{D}$ of the Skorokhod space $D[0,T]$ is equal to the sigma-field generated by finite-dimensional cylinders. Hence, we have that $F\in\sigma(X_{t},0\leq t\leq T)\subset\mathcal{F}^{X}_{T}$ and $G\in\sigma(Y_{t},0\leq t\leq T)\subset\mathcal{F}^{Y}_{T}$. By Theorem 4.2, we find square-integrable predictable processes $\varphi^{X}$, $\psi^{X}$, $\varphi^{Y}$ and $\psi^{Y}$ such that $F=E(F)+F_{B}+F_{N}$ and $G=E(G)+G_{B}+G_{N}$ with probability $1$, where
@@
We only provide the proof for $E(F_{B}G_{N})=0$, as the other equations can be proved similarly. We define a martingale-valued measure $M$ as follows: for $t\geq 0$ and $A\subset\mathbb{R}^{2}$, we define
@@
Let $\mu(t,A)=E(M(t,A)^{2})$. By independence between $\widetilde{N}$ and $B^{X}$, we get that
@@
Let $f(t,x,y)=1_{x=y=0}\varphi^{X}(t)$ and $g(t,x,y)=1_{y\neq 0}\psi^{Y}(t,y)$. Then $F_{B}$ is equal to the stochastic integration $\int_{0}^{T}f(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ and $G_{N}$ is equal to the stochastic integration $\int_{0}^{T}g(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$. For the general theory of stochastic integration against a certain type of martingale-valued measure, we refer to [App09, Chapter 4]. By Itô’s isometry for stochastic integrals (see [App09, Theorem 4.2.3 and Exercise 4.2.4]), we have that
@@
$\left\{\begin{aligned} \operatorname{Var}(F)&=E(F_{B}^{2})+E(F_{N}^{2}),\\ \operatorname{Var}(G)&=E(G_{B}^{2})+E(G_{N}^{2}),\\ \operatorname{Cov}(F,G)&=E(F_{B}G_{B})+E(F_{N}G_{N}).\end{aligned}\right$. (19)
@@
By the Cauchy-Schwarz inequality, we obtain that
@@
Integrating both sides of (22) against $\mathrm{d}t$ and using the Cauchy-Schwarz inequality, we prove (21). Finally, by (19), (20), (21) and the Cauchy-Schwarz inequality, we have that
@@
As $T\to\infty$, $((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})$ converges to $((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. By Lemma 2.3 (lower-semi continuity), we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\liminf_{T\to\infty}R((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})\leq\max(|\rho|,\mathrm{Op}(\nu)),$
@@
In this subsection, we will prove Theorem 1.3. Firstly, we briefly explain the ideas. We have found the maximal correlation coefficient for a two-dimensional Lévy process in Theorem 1.2. Then we apply Theorem 1.2 to a two-dimensional $\alpha$-stable process $(X_{t},Y_{t})_{t\geq 0}$ with $0<\alpha<2$. The Brownian part vanishes. Hence, it suffices to calculate $\mathrm{Op}(\nu)$. For a stable process, $\nu$ takes the special form (14). If $\tau(\mathrm{d}\theta)$ has a density $\tau(\theta)$ with respect to the Lebesgue measure $\mathrm{d}\theta$, the Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$ also has a density $\nu(x,y)$. Moreover, if $\tau(-\theta)=\tau(\theta)$, we have $\nu(-x,-y)=\nu(x,y)$. In this case, we can define a homogeneous kernel
$K(x,y)=\frac{\nu(x,y)}{\sqrt{\nu_{X}(x)\nu_{Y}(y)}}$
of degree 1-1, where
$\nu_{X}(x)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}y,\nu_{Y}(y)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}x,\text{ and }K(\lambda x,\lambda y)=|\lambda|^{-1}K(x,y)$
for all $x,y\in\mathbb{R}$ and $\lambda\neq 0$. The kernel $K(x,y)$ induces a linear operator $K$ on the $L^{2}$-space by
$K\psi(x):=\int_{\mathbb{R}}K(x,y)\psi(y)\,\mathrm{d}y.$
Then the constant $\mathrm{Op}(\nu)$ is just the operator norm
$\|K\|=\sup_{\|\psi\|_{2}>0}\frac{\|K\psi\|_{2}}{\|\psi\|_{2}}.$
@@
Here, $\mathrm{Op}$” is short for the word “operator”. The determination of the norm $\|K\|$ of the homogeneous kernel $K$ of degree 1-1 is a classical problem under the name “Hilbert-Hardy inequality”, see [HLP88, Theorem 319] and [Yan12, Theorem 42.9]. However, in our case, $\tau(\mathrm{d}\theta)$ is not necessarily absolutely continuous with respect to the Lebesgue measure. Moreover, even if it is possible to define the kernel $K(x,y)$ in certain cases, because of the absence of the symmetry $\tau(-\theta)=\tau(\theta)$, $K(x,y)$ is not homogeneous. Instead, $K(x,y)$ is only positively homogeneous, that is, $K(\lambda x,\lambda y)=\lambda^{-1}K(x,y)$ for $x,y\in\mathbb{R}$ and $\lambda>0$. The difference between homogeneity and positive homogeneity results in different expressions for $\|K\|$. Indeed, in the case where $K$ is positive homogeneous, $\|K\|=\mathrm{Op}(\nu)$ is expressed as the spectral norm of a $2\times 2$ matrix, see Lemma 4.4. However, in the case where $K(x,y)$ is a homogeneous kernel, by [Yan12, Theorem 42.9], we have that
$\|K\|=\int_{-\infty}^{\infty}k(x,1)|x|^{-\frac{1}{2}}\,\mathrm{d}x.$
Equivalently, in terms of $\alpha$ and $\tau$, we have that
$\|K\|=\frac{\int_{0}^{2\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\tau(\theta)\,\mathrm{d}\theta}{\sqrt{\int_{0}^{2\pi}|\cos\theta|^{\alpha}\tau(\theta)\,\mathrm{d}\theta\int_{0}^{2\pi}|\sin\theta|^{\alpha}\tau(\theta)\,\mathrm{d}\theta}}.$
We are not aware of existing results for $\mathrm{Op}(\nu)$ in general. Hence, we calculate $\mathrm{Op}(\nu)$ in Lemma 4.4 by an adaptation of the argument leading to the Hilbert-Hardy inequalities. In this way, we subsequently find the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for a two-dimensional stable process. As $\alpha$-stable distributions are marginal distributions of $\alpha$-stable processes, we have the upper bound $R(X_{1},Y_{1})\leq R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. The reverse inequality is a consequence of the convergence of stable random walks towards stable processes, Theorem 1.1 and Lemma 2.3. However, we are unable to generalize Theorem 1.3 to general infinitely divisible distributions. Indeed, it is possible to have $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for a general two-dimensional Lévy process, see Examples 3 and 4.
@@
Then we have that
$\mathrm{Op}(\nu)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&\quad C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&\quad C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $0/0=0$, where $\|\cdot\|_{2}$ denotes the spectral norm.
@@
using the Cauchy-Schwarz inequality, we obtain that
@@
$\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{x}_{+}F_{+}+D^{x}_{-}F_{-}$ (25)
@@
$\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{y}_{+}G_{+}+D^{y}_{-}G_{-}.$ (26)
@@
$A=\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&\quad C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&\quad C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}$
with the convention that $0/0=0$. Let $\|A\|_{2}$ be the spectral norm of $A$. Then by (24), (25) and (26), we have that
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\leq\|A\|_{2}\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}.$
Hence, $\mathrm{Op}(\nu)\leq\|A\|_{2}$. It remains to prove the reverse inequality $\mathrm{Op}(\nu)\geq\|A\|_{2}$. Without loss of generality, we may assume that $\|A\|_{2}>0$. Equivalently, we assume that $D^{x}_{+}+D^{x}_{-}>0$ and $D^{y}_{+}+D^{y}_{-}>0$. For this purpose, for $\varepsilon>0$, we set
@@
where the positive constants $b^{\varphi}_{+}$, $b^{\varphi}_{-}$, $b^{\psi}_{+}$ and $b^{\psi}_{-}$ will be chosen later. Then we have that
@@
$(\sqrt{D^{x}_{+}}b^{\varphi}_{+},\sqrt{D^{x}_{-}}b^{\varphi}_{-})A\begin{pmatrix}\sqrt{D^{y}_{+}}b^{\psi}_{+}\\ \sqrt{D^{y}_{-}}b^{\psi}_{-}\end{pmatrix}\\ =\|A\|_{2}\sqrt{(D^{x}_{+}(b^{\varphi}_{+})^{2}+D^{x}_{-}(b^{\varphi}_{-})^{2})(D^{y}_{+}(b^{\psi}_{+})^{2}+D^{y}_{-}(b^{\psi}_{-})^{2})}.$
Then we get that
$\mathrm{Op}(\nu)\geq\lim_{\varepsilon\to 0}\frac{\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)}{\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}}=\|A\|_{2}.$
Currently, we have found an expression for $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. Clearly, we have that
@@
On the other hand, consider the random walk $(S_{n},T_{n})_{n\geq 0}$ such that its increment $(S_{n+1}-S_{n},T_{n+1}-T_{n})$ has the same distribution as $(X_{1},Y_{1})$. By Theorem 1.1, we have that